a daughters age is 32 years younger than her farther. In 4 years, the farther will be 5 times as old as hid daughter. How old is each now?
Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39620)

d = daughter's age. f = father's age.

d=f-32 and f+4=5(d+4)

Answer by stanbon(75887)

a daughters age is 32 years younger than her father. In 4 years, the father will be 5 times as old as his daughter. How old is each now?
-----
Equations:
d = f-32
(f+4) = 5(d+4)
-----
Substitute for "d" and solve for "f":
f+4 = 5(f-32+4)
f+4 = 5(f-28)
f+4 = 5f-140
4f = 144
father's age = 144/4 = 36
-----
Solve for "d":
daughter's age = 36-32 = 4
============
cheers,
Stan H.
==============